What is the equation in slope-intercept form of a line containing the points (1/2,6) and (-1/2,8)?

Answer by stanbon(75887)

What is the equation in slope-intercept form of a line containing the points (1/2,6) and (-1/2,8)?
------------------------
slope = (8-6)/(-1/2-1/2) = 2/-1 = -2
------
Form: y = mx+b
6 = -2(1/2) + b
----------
6 = -1 + b
b = 7
-----------------
Equation:
y = -2x + 7
==============
Cheers,
Stan H.
==============